a cell contains 180m3 of gas at 7400Pa and a machine which is turned on and expands. During this process the machine gives off 260kJ of heat to the gas and the internal engery is determined to be -69kJ. What is the final volume of the cell assuming pressure stays constant
1
Expert's answer
2020-04-23T13:06:05-0400

Isobaric process: W = p∆V

∆U = Q - W - iternal energy change

∆U = Q - p∆V

∆V = (Q - ∆U) / p

∆V = (260000 J + 69000 J) / 7400 Pa

∆V = 35 m3

Final volume: = 180 + 35 = 215 m3